Dive Number: 3954
December 13, 2003
Port Observer: Brian Glazer
Starboard Observer: Frank Stewart
Onboard echem & Sipper functional; AIS temp malfunction.
We splashed at 13:59 GMT, descended at 14:02 @ 30m/min, and saw a school of yellowfin tuna at 25 m. Poor navigation until 550+ m. Reached bottom at 15:28 GMT (x4533, y77966 @ 2502 m).
Gained neutral buoyancy and arrived at P vent at 15:33 GMT. Most Alvinellids on P-middle, but are very thin, buried right on sulfide structure and Pat made the call to move on to Bio 9 for Alvinellid collection.
Viewed newly forming Alvinellid carpet en route to Bio 9 (x4608y, 77976, 15:39 GMT) with intention to return for scanning if time permitted after Alvinellid sampling.
Arrived at Bio 9 at 15:40 GMT (x4613, y77981, 2508 m, HDG037). Began temperature & echem scanning (scans 13–177) of 10 tubes. Collected Sippers from 7 tubes.
Sipper 1: 15:52 GMT @ 41°C
Began grabbing worms at 16:42 GMT. Collected 8 into ARTIE Chamber 1 & confirmed RNA Later venting at 17:06 GMT.
Set up on SE corner of Alvinellid carpet (x4608, y77974, 2512 m, HDG 308, 17:08 GMT). Scanned echem (scans 180–187) & temperatures ranging from 50–70°, with a hot spot of 240°C.
Transit past Mr. Nuzzio on way to Tica. Arrived at southern end of Tica at 17:26 GMT (x2512, y78151, 2512 m, HDG080) for Riftia collection. Temperatures were near ambient, with one spot of 6°C. Echem scans 190–196. 2 Riftia into RNA Later chamber 3 & verified RNA Later at 17:44 GMT.
Arrived at Tica beehive scars at 17:48 GMT (x4571, y78170, 2512 m, HDG187) and started setting up for deployment of ASU frying pan. Orifice temperatures ranged from 188° @ 10 cm above to 322°C at the orifice. Positioned frying pan over lower orifice with some overlap to upper orifice at 18:01 GMT. Fired Sipper 8 (18:03 GMT @ 52°C) in venting on lower side of pan & Sipper 9 (18:07 GMT @ 13°C) from just above pan in venting waters. Echem scans 220–237.
Deployed Fries #53 in Riftia clump at 18:13 GMT, and began collecting others (39, 40, 41, 42) at 18:15 GMT. Fries #39 had fallen out of Riftia clump. Collected final two Fries (45 & 46) in periphery at 18:29 GMT (x4604, y78152, 2502 m, HDG067).
Check up on Insect at 18:40 GMT. Both Riftia Remoras had fallen out of tubes, but arms were still positioned well. Only able to temperature/echem scan red-taped Vimco arm due to sub & Insect positioning. Unable to redeploy Remoras for same reason. Temps ranged from 4.5 to 9°C. Took echem scans 247–253.
Transit to Biovent 18:42 GMT. Arrived & sighted ASU frying pan at Biovent at 19:38 GMT (x4334, y79174, 2505 m, HDG017). Sipper #10 taken @ 99°C in fluid venting through ASU frying pan. Sipper #11 taken from back side @ 20°C. Echem scans 264–280. Recovered pan (19:48 GMT) and scanned orifice. Temperatures ranged up to 100°C within 5 cm above orifice. Echem scans 313–325. Sipper 12 taken @ 100°C. Echem scans 328–333 taken at 120°C at 4000 mVs-1.
Dropped weights at 20:00 GMT.
|Copyright University of Delaware, November 2003|